# An operational approach to classifying measurement incompatibility

### v1

@@ IV.1 Coarse-graining of outcomes
Table 3: The incompatibility status for all possible coarse-graining (CG) of outcomes for Alice’s measurements ($A_{1}$ and $A_{2}$) as inferred from the statistics of CGLMP experiment. Here “yes” denotes that under the particular CG of $A_{1}$ and $A_{2}$, the measurement remains incompatible. On the other hand, “????” denotes that we can not make any conclusion about their incompatibility.
IV.2 Convex mixing of measurements
@@
Here, one can simply use the result that any two binary-outcome incompatible measurements violate the Bell-CHSH inequality Wolf et al. (2009). Considering the convex mixture of measurements for any three binary-outcome measurements, this result can be used to witness full incompatibility w.r.t. convex mixing of measurements.
@@
For three or more numbers of measurements we can witness whether the measurements are fully-incompatible w.r.t convex-mixture in a semi-device independent manner by constructing a suitable random access codes task.
@@
Three noisy Pauli measurements of Eq.(16) are witnessed to be fully incompatible w.r.t convex-mixing via RAC iff: $\sqrt{\frac{2}{3}}<\nu\leq 1.$
@@
Let us denote the measurements of Eq.(16) as $\mathcal{M}_{0}=\frac{1}{2}\left(\mathbbm{1}+\nu_{0}\sigma_{x}\right)$, $\mathcal{M}_{1}=\frac{1}{2}\left(\mathbbm{1}-\nu_{0}\sigma_{x}\right)$, $\mathcal{N}_{0}=\frac{1}{2}\left(\mathbbm{1}+\nu_{1}\sigma_{y}\right)$, $\mathcal{N}_{1}=\frac{1}{2}\left(\mathbbm{1}-\nu_{1}\sigma_{y}\right)$, $\mathcal{R}_{0}=\frac{1}{2}\left(\mathbbm{1}+\nu_{2}\sigma_{z}\right)$, and $\mathcal{R}_{1}=\frac{1}{2}\left(\mathbbm{1}-\nu_{2}\sigma_{z}\right)$, and consider the measurements $Q^{i}$ with $i\in\{1,2\}$, to be formed by the convex-mixing of $\mathcal{M}$ and $\mathcal{N}$ as
@@
$Q^{2}=\{p\,\,M_{0}+(1-p)N_{1},p\,\,M_{1}+(1-p)N_{0}\}.$
In order to witness the incompatibility status of $(Q^{i},R)$, one can construct a (2,2,2) RAC task. Now, when Bob has to guess Alice’s 1st bit, he performs the measurement $Q^{i}$ which can be realized by the convex mixing of measurements $\mathcal{M}$ and $\mathcal{N}$ as follows: Bob tosses a biased coin which gives a ‘Head’ with probability $p$ and a ‘Tail’ with probability $(1-p)$. On obtaining ‘Head’, Bob performs measurement $\mathcal{M}$, and for ‘Tail’ he performs measurement $\mathcal{N}$. For the second bit, he performs measurement $\mathcal{R}$. The maximum average success probability $P(2,2,2)$ that can be obtained in the RAC task (considering $\nu_{0}=\nu_{1}=\nu_{2}$) is:
@@
It can be shown that for all possible convex-mixing the maximum average success probability is the same as Eq.(62). Now, to be fully incompatible w.r.t convex-mixing $P(2,2,2)>\frac{3}{4},\,\,\text{which implies}\,\,\nu^{2}(p^{2}-p+1)>\frac{1}{2}$ (by Eq.(62)). Since the minimum value of $(p^{2}-p+1)$ is $\frac{3}{4},$ so for $\nu>\sqrt{\frac{2}{3}}$ the measurements are fully incompatible w.r.t convex-mixing.
In Observation 4, we have found that this kind of three qubit measurements with noise become fully incompatible w.r.t convex mixture for $\nu>\sqrt{\frac{2}{3}}$. So, we can conclude that when these three measurements are fully incompatible w.r.t convex mixing, quantum advantage is obtained in a suitable Random Access Code task.
@@
Examples.— Consider the following three rank-one projective measurements acting on $\mathbb{C}^{3}$:
@@
and $Z=\{|0_{z}\rangle\langle 0_{z}|,|1_{z}\rangle\langle 1_{z}|,|2_{z}\rangle\langle 2_{z}|\},$ (63)
@@
We take the following convex mixture of the measurements $X$ and $Y$: $A=\{p|0_{x}\rangle\langle 0_{x}|+(1-p)|0_{y}\rangle\langle 0_{y}|,p|1_{x}\rangle\langle 1_{x}|+(1-p)|1_{y}\rangle\langle 1_{y}|,p|2_{x}\rangle\langle 2_{x}|+(1-p)|2_{y}\rangle\langle 2_{y}|\}$ with $0\leq p\leq 1$. We now consider the $(2,3,3)$ RAC game involving the two measurements $A$ and $Z$. The maximum average success probability $P(2,3,3)$ for this RAC task is given by Carmeli et al. (2020),
@@
Next, consider the following measurement (taking an arbitrary convex mixture of $X$ and $Y$) $A_{i,j,k,l,m,n}=\{p|i_{x}\rangle\langle i_{x}|+(1-p)|j_{y}\rangle\langle j_{y}|,p|k_{x}\rangle\langle k_{x}|+(1-p)|l_{y}\rangle\langle l_{y}|,p|m_{x}\rangle\langle m_{x}|+(1-p)|n_{y}\rangle\langle n_{y}|\}$ with $0\leq p\leq 1$, $i,j,k,l,m,n\in\{0,1,2\}$, $i\neq k$, $k\neq m$, $m\neq i$ and $j\neq l$, $l\neq n$, $n\neq j$. Note that the aforementioned measurement $A$ is nothing but $A_{0,0,1,1,2,2}$ in the present notation. Following a similar calculation, it can be shown that any such $A_{i,j,k,l,m,n}$ is incompatible with $Z$ for all $p\in[0,1]$.
@@
A generic description for the semi-device independent witness of incompatibility.— Consider a measurement assemblage of $n$ measurements: $M=\{M_{0},M_{1},\ldots,M_{n-1}\}.$ We make $m$ partitions $S=\{S_{i}\}_{i=0}^{m-1}$, $\bigcup_{i}S_{i}=M,\,\text{and}\,S_{i}\bigcap S_{j}=\phi,\,\,\forall\,\,i,j\,\,\text{with}\,\,i\neq j.$ Our purpose is to operationally witness the incompatibility of $m$ measurements produced from the convex mixing of $n$ measurements from each of the $m$ partitions. We can construct a $(m,d,d)$ RAC game where Alice has an $m$ dit input message, viz., $x=(x_{0},x_{1},\ldots,x_{m-1})$. Depending upon the message, she encodes it in a qudit and sends it to Bob, who on the other hand, gets input $y\in\{0,1,\ldots,m-1\}$ and accordingly, he has to predict the value of the corresponding bit $x_{y}$. He performs the measurement which is obtained by the convex mixtures of the measurements from the partition $S_{y}$ and declares the outcome of the measurement. Now, if the success probability $P(m,d,d)$ is greater than $P_{CB}(m,d,d)$ Saha et al. (2023) for all possible convex mixtures, we can operationally detect that the $m$ partitioned measurements are incompatible under convex mixing.
V Conclusions
@@
Measurement incompatibility is the spectacle in quantum theory that a set of measurements cannot be performed jointly on arbitrary systems. It is one of the fundamental ingredients for non-classical correlations and merits of quantum information science. Incompatibility offers a complex structure with different layers as the number of measurements and the dimension on which the measurements act increases. Thus, understanding the different levels of incompatibility with respect to elementary classical operations is of paramount importance both from the foundational and practical perspectives. In this work, we have considered the two most general classical operations viz., coarse-graining of different outcomes of measurements, and convex-mixing of different measurements.
@@
Through our present analysis we have investigated the different levels of incompatibility arising under the above classical operations. Since environmental effects are ubiquitous in practical scenarios, here we have investigated the tolerance thresholds for maintaining measurement incompatibility against noise under classical operations. Furthermore, we have developed a method to operationally witness different levels of measurement incompatibility in the device-independent framework involving Bell-type experiments, and also in the semi-device-independent framework involving prepare-and-measure experiments.
@@
Several examples have been provided to illustrate the efficacy of the operational witnesses proposed here. Our results are particularly useful for the purpose of comparing measurements in terms of their degree of incompatibility. For example, measurements that remain fully incompatible with respect to coarse-graining of outcomes (or convex-mixing of measurements) show stronger incompatibility compared to those that are not fully incompatible, thus facilitating legitimate choice of measurements in an information-processing task where a high degree of incompatibility is required.
@@
Our present study motivates future work on several open directions. The condition for full incompatibility for two projective measurements with respect to coarse-graining of outcomes can be extended for more number of projective measurements. Similarly, the criterion for full incompatibility of projective measurements with respect to convex mixing could be generalized for any set of measurements. Further investigation would be required to see if device-independent witnesses could be formulated for more than three outcome measurements with respect to coarse-graining and more than three measurements with repect to convex mixing. It will be interesting to look for examples where the full incompatibility with respect to coarse-graining can be inferred in a device-independent way from a single experimental statistics.
VI Acknowledgements

### v2

@@ V.1 Coarse-graining of outcomes
Table 3: Incompatibility status for all possible coarse-graining (CG) choices of outcomes of Alice’s measurements ($A_{1}$ and $A_{2}$) in the CGLMP scenario. Here $\Delta S_{T}$ refers to the amount of CH-inequality violation obtained in the theory and $\Delta S_{E}$ refers to the CH-inequality violation calculated from the states and measurements viz. (42) and (43) realised in the experiment Zhang et al. (2024). Here “yes” denotes that under the particular CG of $A_{1}$ and $A_{2}$, the measurements can be witnessed to be incompatible as they give CH-inequality violation. On the other hand, “????” denotes that we can not make any conclusion about their incompatibility under those particular CG as there is no violation of CH-inequality in those cases.
Experimental realization.— Recently, an experimental study on CGLMP inequality has been performed Zhang et al. (2024) where the two parties share orbital angular momentum entanglement in a scenario of multiple settings and outcomes. In their experiment, for $d=3$ they prepared an orbital angular momentum entangled state which is in $\mathbb{C}^{3}\otimes\mathbb{C}^{3}$ of the form:
$|\psi\rangle$ $=$ $0.596|+1\rangle_{A}|-1\rangle_{B}+0.529|+2\rangle_{A}|-2\rangle_{B}$ (42)
$+0.604|-1\rangle_{A}|+1\rangle_{B},$
with $|+l\rangle_{A}$ corresponds to the orbital angular momentum eigenstate of the signal photon (A) with orbital angular momentum $+l\hbar\,\,\text{and}\,\,|-l\rangle_{B}$ corresponds the orbital angular momentum eigenstate of the corresponding idler photon (B) with orbital angular momentum $-l\hbar.$ The von Neuman measurements of Alice and Bob are given by $A_{i}\equiv\{|\Gamma^{i}_{s}\rangle_{A}\langle\Gamma^{i}_{s}|\}\,\,\text{and}\,\,B_{j}\equiv\{|\Theta^{j}_{t}\rangle_{B}\langle\Theta^{j}_{t}|\}$ respectively, where
$|\Gamma^{i}_{s}\rangle_{A}=\frac{1}{\sqrt{3}}(|+1\rangle_{A}+\omega^{s+\sigma_{i}}|+2\rangle_{A}+\omega^{2(s+\sigma_{i})}|-1\rangle_{A}),$ (43)
and
$|\Theta^{j}_{t}\rangle_{B}=\frac{1}{\sqrt{3}}(|-1\rangle_{B}+\omega^{-t-\gamma_{j}}|-2\rangle_{B}+\omega^{2(-t-\gamma_{j})}|+1\rangle_{B}),$ (44)
with $i,j\in\{1,2\}\,\,\text{and}\,\,s,t\in\{0,1,2\},\sigma_{1}=\frac{1}{4},\sigma_{2}=\frac{3}{4},\gamma_{1}=\frac{1}{2}\,\,\text{and}\,\,\gamma_{2}=0,\,\,\omega=\text{exp}(\mathbbm{1}\frac{2\pi}{3}).$ For $d=3$ the incompatibility status of the measurements (see (43)) used in their experimental arrangements before and after coarse-graining is depicted in TABLE 3.
Semi-device-independent witness
In the semi-device-independent approach in prepare-and-measure experiments, we do not have any prior knowledge of the internal functioning of the measurement device; however, we assume the dimension of the system on which the measurements act. To witness different levels of incompatibility, here we focus on a class of communication tasks, namely, $(2,\overline{d},d)-$ random access code tasks (RAC). In this task, the sender, Alice, gets two-dit string input message $(x_{1},x_{2})$ with $x_{1},x_{2}\in[\overline{d}],$ and can communicate a $d-$dimensional system to the receiver, Bob, who wants to guess the value of any of the two dits, i.e. $x_{1}$ or $x_{2}$, randomly. The notation of $(2,\overline{d},d)-$RAC is adopted from Saha et al. (2023). If any two POVMs, each with $\overline{d}$ outcomes acting on $\mathbb{C}^{d}$ are jointly measurable, the average success probability of this task,
$P(2,\overline{d},d)\leq P_{CB}(2,\overline{d},d)=\frac{1}{2}\left(1+\frac{d}{\overline{d}^{2}}\right),$ (45)
where $P_{CB}(2,\overline{d},d)$ is an upper bound on average success probability using two compatible measurements Carmeli et al. (2020); Saha et al. (2023). It turns out any two incompatible rank-one projective measurements can be witnessed through $(2,d,d)-$RAC Saha et al. (2020). We prove the following result for full incompatibility of two 3-outcome rank-one projective measurements w.r.t. coarse-graining.
Theorem 4.
Two 3-outcome rank-one projective measurements, $M=\left\{|\phi_{0}\rangle,|\phi_{1}\rangle,|\phi_{2}\rangle\right\}$ and $N=\left\{|\psi_{0}\rangle,|\psi_{1}\rangle,|\psi_{2}\rangle\right\}$, can be witnessed to be fully incompatible w.r.t. coarse-graining via RAC if and only if $20<|\langle\phi_{i}|\psi_{j}\rangle|<\frac{4}{5},\quad\forall\,\,i,j=0,1,$ 2 .
Proof.
$M=\left\{|0\rangle\langle 0|,|1\rangle\langle 1|,|2\rangle\langle 2|\right\}\equiv\left\{|\phi_{i}\rangle\langle\phi_{i}|\right\},$
$N=\left\{|\psi_{1}\rangle\langle\psi_{1}|,|\psi_{2}\rangle\langle\psi_{2}|,|\psi_{3}\rangle\langle\psi_{3}|\right\},$ (46)
Let us first consider the coarse-graining of the second and third outcomes for both measurements. So, the new measurement pair becomes:
This pair of measurements appears in $(2,2,3)-$RAC and the compatibility bound in this case is,
$P_{CB}(2,2,3)=\frac{1}{2}\left(1+\frac{3}{2^{2}}\right)=\frac{7}{8}$ (48)
by Eq.(45) and this bound is tight. $M^{(2,3)}\,\,\text{and}\,\,N^{(2,3)}$ will give advantage in $(2,2,3)-$RAC if
$\sum_{x,y=0}^{1}||M^{(2,3)}_{x}+N^{(2,3)}_{y}||=\sum_{i=1}^{4}||A_{i}||>7,$ (49)
with
$A_{1}=M^{(2,3)}_{0}+N^{(2,3)}_{0}=|0\rangle\langle 0|+|\psi_{1}\rangle\langle\psi_{1}|,$
$A_{2}=M^{(2,3)}_{0}+N^{(2,3)}_{1}=\mathbbm{1}+|0\rangle\langle 0|-|\psi_{1}\rangle\langle\psi_{1}|,$
$A_{3}=M^{(2,3)}_{1}+N^{(2,3)}_{0}=\mathbbm{1}-|0\rangle\langle 0|+|\psi_{1}\rangle\langle\psi_{1}|,$
$A_{4}=M^{(2,3)}_{1}+N^{(2,3)}_{1}=2\mathbbm{1}-|0\rangle\langle 0|-|\psi_{1}\rangle\langle\psi_{1}|,$ (50)
and ||.||||.||| | . | | denotes the maximum eigenvalue of an operator. Now, $|\psi_{1}\rangle$ can be written as follows
$\langle 0|u\rangle=0\,\,\text{and}\,\,|\langle 0|\psi_{1}\rangle|^{2}+|\langle u|\psi_{1}\rangle|^{2}=1,$
for some vector $|u\rangle.$ So, $A_{1}$ can be written as a $(2\times 2)$ matrix in the basis $\{|0\rangle,|u\rangle\},$
$A_{1}=\begin{pmatrix}1+|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (52)
the maximum eigenvalue of $A_{1}\,\,\text{is}\,\,1+|\langle 0|\psi_{1}\rangle|.$ Thus,
$||A_{1}||=1+|\langle 0|\psi_{1}\rangle|.$ (53)
Similarly, $A_{2}$ can be expressed in a block diagonal matrix in the ortho-normal basis $\{|0\rangle,|u\rangle,|v\rangle\},$
$A_{2}=\begin{pmatrix}\Gamma&0\\ 0&1\end{pmatrix}\,\,\text{where}\,\,\Gamma=\begin{pmatrix}2+|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&1+|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (54)
and
$||A_{2}||=2+|\langle 0|\psi_{1}\rangle|.$ (55)
$A_{3}$ and $A_{4}$ can also be expressed in a block diagonal matrix in the basis $\{|0\rangle,|u\rangle,|v\rangle\},$
$A_{3}=\begin{pmatrix}\Sigma&0\\ 0&1\end{pmatrix},\ \Sigma=\begin{pmatrix}|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&1+|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (56)
and
$||A_{3}||=1+\sqrt{1-|\langle 0|\psi_{1}\rangle|^{2}}.$ (57)
$A_{4}=\begin{pmatrix}\Xi&0\\ 0&2\end{pmatrix},\ \Xi=\begin{pmatrix}1-|\langle 0|\psi_{1}\rangle|^{2}&-\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ -\langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&2-|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (58)
and $||A_{4}||=2.$
[…]
V.2 Disjoint-convex-mixing of measurements
We now study the operational witness of incompatibility w.r.t. disjoint-convex-mixing of measurements.
@@ V.2 Disjoint-convex-mixing of measurements
Given a set of $n$ measurements, we can make $k$ partitions of it, where $k\in$$\{2,\cdots,n\}$. If the $k$ number of measurements obtained by the disjoint-convex-mixing of the measurements provides a violation of Bell inequalities for all possible disjoint-convex-mixing and permutations, then the measurements are witnessed to be $k-$incompatible w.r.t. disjoint-convex-mixing in a device-independent way. Using the result that any two binary-outcome incompatible measurements violate the Bell-CHSH inequality Wolf et al. (2009a), we can witness 2-incompatibility w.r.t. disjoint-convex-mixing from any set of binary-outcome measurements.
@@
For three or more numbers of measurements, we can witness whether the measurements are fully-incompatible w.r.t. disjoint-convex-mixture in a semi-device independent manner by constructing suitable random access code tasks.
@@
Three noisy Pauli measurements of Eq. (3) with equal noise ($\nu=\nu_{0}=\nu_{1}=\nu_{2}$) are witnessed to be fully incompatible w.r.t. disjoint-convex-mixing via RAC if and only if $\sqrt{2/3}<\nu\leqslant 1.$
@@
Consider the measurements $\{Q^{i}\}$ with $i\in\{1,2\}$ for two possible permutations of outcomes, to be formed by the disjoint-convex-mixing of $\{M\}$ and $\{N\}$ as
@@
$Q^{2}=\{p\,\,M_{0}+(1-p)N_{1},p\,\,M_{1}+(1-p)N_{0}\}.$ (59)
To witness the incompatibility status of $(Q^{i},R)$, one can construct a $(2,2,2)-$RAC task. Now, when Bob has to guess Alice’s 1st bit, he performs the measurement $\{Q^{i}\}$ defined by (V.2), which is realized by the disjoint-convex-mixing of measurements $\{M\}$ and $\{N\}$. For the second bit, he performs the measurement $\{R\}$. The maximum average success probability $P(2,2,2)$ that can be obtained in the RAC task is:
@@
It can be shown that for all possible disjoint-convex-mixing, the maximum average success probability is the same as Eq. (60). Now, to be fully incompatible w.r.t. disjoint-convex-mixing $P(2,2,2)>3/4,\,\,\text{which implies}\,\,\nu^{2}(p^{2}-p+1)>\frac{1}{2}$ (by Eq.(60)). Since the minimum value of $(p^{2}-p+1)$ is $\frac{3}{4},$ so for $\nu>\sqrt{2/3}$ the measurements are fully incompatible w.r.t. disjoint-convex-mixing.
@@
Let us consider another example of three rank-one projective measurements acting on $\mathbb{C}^{3}$:
@@
$Z=\{|0_{z}\rangle\langle 0_{z}|,|1_{z}\rangle\langle 1_{z}|,|2_{z}\rangle\langle 2_{z}|\},$ (61)
@@
We take the following disjoint-convex-mixing of the measurements $X$ and $Y$: $A=\{p|0_{x}\rangle\langle 0_{x}|+(1-p)|0_{y}\rangle\langle 0_{y}|,p|1_{x}\rangle\langle 1_{x}|+(1-p)|1_{y}\rangle\langle 1_{y}|,p|2_{x}\rangle\langle 2_{x}|+(1-p)|2_{y}\rangle\langle 2_{y}|\}$ with $0\leq p\leq 1$. We now consider the $(2,3,3)$ RAC game involving the two measurements $A$ and $Z$. The maximum average success probability $P(2,3,3)$ for this RAC task is given by,
@@
Next, consider the following measurement (taking an arbitrary disjoint-convex-mixing of $X$ and $Y$) $A_{i,j,k,l,m,n}=\{p|i_{x}\rangle\langle i_{x}|+(1-p)|j_{y}\rangle\langle j_{y}|,p|k_{x}\rangle\langle k_{x}|+(1-p)|l_{y}\rangle\langle l_{y}|,p|m_{x}\rangle\langle m_{x}|+(1-p)|n_{y}\rangle\langle n_{y}|\}$ with $0\leq p\leq 1$, $i,j,k,l,m,n\in\{0,1,2\}$, $i\neq k$, $k\neq m$, $m\neq i$ and $j\neq l$, $l\neq n$, $n\neq j$. Note that the aforementioned measurement $A$ is denoted by $A_{0,0,1,1,2,2}$ following the present notation. Following a similar calculation, it can be shown that any such $A_{i,j,k,l,m,n}$ is incompatible with $Z$ for all $p\in[0,1]$.
@@
A generic description for the semi-device independent witness of incompatibility.— Consider a measurement assemblage of $n$ measurements having $\overline{d}$ outcomes, =$\mathcal{M}=\{M_{z|x}\}_{z,x}$ where $x\in[n],z\in[\overline{d}]$, and the measurements act on $\mathbbm{C}^{d}$. We make $k$ partitions: $S=\{S_{i}\}_{i=0}^{k-1}$, $\bigcup_{i}S_{i}=\mathcal{M},\,\text{and}\,S_{i}\bigcap S_{j}=\phi,\,\,\forall\,\,i,j\,\,\text{with}\,\,i\neq j.$ Our purpose is to operationally witness the incompatibility of $k$ measurements produced from the disjoint-convex-mixing of $n$ measurements from each of the $k$ partitions. We can construct a $(k,\overline{d},d)-$RAC where Alice has an $k$ dit input message, viz., $x=(x_{0},x_{1},\ldots,x_{k-1})$. Depending upon the message, she encodes it in a qudit and sends it to Bob, who on the other hand, gets input $y\in[k]$ and accordingly, he has to predict the value of the corresponding bit $x_{y}$. He performs the measurement, which is obtained by the disjoint-convex-mixing of the measurements from the partition $S_{y}$ and declares the outcome of the measurement. Now, if the success probability $P(k,\overline{d},d)$ is greater than $P_{CB}(k,\overline{d},d)$ Saha et al. (2023) for all possible disjoint-convex-mixing and permutations of measurement outcomes, we can operationally witness $k-$incompatibility under disjoint-convex-mixing.
VI Conclusions
@@
Measurement incompatibility is a feature in quantum theory that a set of measurements cannot be performed jointly on arbitrary systems Gühne et al. (2023). It is one of the fundamental ingredients for non-classical correlations and merits of quantum information science. Incompatibility offers a complex structure with different layers as the number of measurements and the dimension on which the measurements act increases Uola et al. (2021); Egelhaaf et al. (2024); Rahimi-Keshari et al. (2021). Thus, understanding the different levels of incompatibility with respect to elementary classical operations is of paramount importance from both the foundational and practical perspectives. In this work, we have considered the two most general classical operations viz., coarse-graining of different outcomes of measurements, and disjoint-convex-mixing of different measurements.
@@
Through our present analysis, we have investigated the different levels of incompatibility arising under the above classical operations. Since environmental effects are ubiquitous in practical scenarios, the tolerance thresholds for maintaining measurement incompatibility against noise under classical operations are investigated here. Furthermore, we have developed a method to operationally witness different levels of measurement incompatibility in the device-independent framework involving Bell-type experiments, and also in the semi-device-independent framework involving prepare-and-measure experiments.
@@
Several examples have been provided to illustrate the efficacy of the operational witnesses proposed here. Our results are particularly useful for the purpose of comparing measurements in terms of their degree of incompatibility. For example, measurements that remain fully incompatible with respect to coarse-graining of outcomes (or disjoint-convex-mixing of measurements) show stronger incompatibility compared to those that are not fully incompatible, thus facilitating the legitimate choice of measurements in an information-processing task where a high degree of incompatibility is required. Our work allows for experimental implementation and one can infer more about these layers of incompatibility by employing these operations in the statistics of previously conducted experiments Zhang et al. (2024). Further, our formalism can be applied to predict the threshold values of system parameters enabling observations of incompatibility layers in future experiments.
@@
Our present study motivates future work in several open directions. The condition for full incompatibility for two projective measurements with respect to the coarse-graining of outcomes can be extended for more projective measurements. Similarly, the criterion for full incompatibility of projective measurements with respect to disjoint-convex-mixing could be generalized for any set of measurements. It will be interesting to look for examples where the full incompatibility with respect to coarse-graining can be inferred in a device-independent way from single experimental statistics.
Acknowledgements